OpenStudy (anonymous):

*Advance Algebra With Financial Applications* [Work Is Shown] Garrett has an annuity that pays $2,460 at the beginning of each year. If the economy grows at a rate of 2.35% semiannually, what is the value of the annuity if he received it in a lump sum now rather than over a period of nine years? $40,168.22 $39,701.72 $20,084.11 $19,850.86 -------- FV = 2460((1+0.0235/2)^(2*9) - 1 )/( 0.0235/2 )

OpenStudy (amistre64):

present value of an annuity due

OpenStudy (amistre64):

how do we interpret: grows at a rate of 2.35% semiannually

OpenStudy (anonymous):

divide or time it by 2 ?

OpenStudy (amistre64):

if you receive money know; and money in a year from now ... the intermediary rates are moot; all that matters is the yearly rates i believe

OpenStudy (amistre64):

2.35 doubled is 4.70 by the next year PV of B now is B PV of B beginning of year is 2.35 before the increase PV of B beginning of 2years is 2.35+4.70 before the increase PV of B beginning of 3years is 2.35+2(4.70) before the increase etc does that make sense?

OpenStudy (amistre64):

\[B/r_0+B/r_1+B/r_2+...+B/r_9=Present~value\]

OpenStudy (amistre64):

excel would be useful for this :)

OpenStudy (anonymous):

but is there an easier formula

OpenStudy (amistre64):

not really; which is why excel is useful. i got no idea how to work a financial calculator with that stuff

OpenStudy (amistre64):

am i interpreting the growth rate correctly?

OpenStudy (anonymous):

was mine ok ? FV = 2460((1+0.0235/2)^(2*9) - 1 )/( 0.0235/2 )

OpenStudy (amistre64):

i cant say .... looks like it has the parts ... but im not sure how youre interpreting the groath rate is a /2

OpenStudy (anonymous):

2.35% semiannually <---- 0.235*2 ?

OpenStudy (amistre64):

it gets bumped up 2 times a year by .0235 ... and since our payments are at the start of the year. that would amount to a Payment of 2460 that is .0470 less than what it was worth a year before

OpenStudy (amistre64):

the first year being .0235 less then that year before

OpenStudy (amistre64):

0: B0 = 2460(1) 1: B1 = B0 (1-.0235) 2: B2 = B1 (1-.0470) 3: B3 = B2 (1-.0470) ... is this right?

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

then the sum of those values i get to about 20430 with an excel run

OpenStudy (amistre64):

i cant be sure of it tho wince i have some doubts as to my interpretation of the information

OpenStudy (anonymous):

do you know the answer ?

OpenStudy (anonymous):

I never have trouble with these questions before .

OpenStudy (amistre64):

im not sure what the answer would be at the moment no what answer do you get?

OpenStudy (anonymous):

20,084.11

OpenStudy (anonymous):

wild guess

OpenStudy (anonymous):

@amistre64 dont waist your time , thank's alot for your time :D . I'm going to head to tutoring later for this lesson

OpenStudy (amistre64):

what troubling me is: Present value is defined as the amount of money you would have to invest at todays interest rates to get the value of the Future amount. We are not given a present day interest rate to calculate this with. 2460 = PV(1+i)^n. PV = 2460/(1+i)^n This does not account for inflation (economic growth rate) the Present value of a sum of Future money adjusted for inflation is similar, but uses the inflation rate: PV = 2460/(1+g)^n Since no current interest rate is available to determine what i think is a valid definition of Present Value .... Does this just imply that the sum totals due to inflation only?